LadyRavyRay:

Find the volume of the prism. Step 1 out of 2: Find the base area B of the trapezoidal prism. B = 1/2 (b1 + b2)h Area of a trapezoid with bases of lengths b1 and b2 and height h = 1/2 (11 + 10)(8) Substitute 11 for b1, 10 for b2, and 8 for h. = cm2

LadyRavyRay:

EWO!

ZombieLover:

Hi! Do you know how to find the solution?

LadyRavyRay:

S-somewhat

ZombieLover:

ok what would be the base, height and other base?

ZombieLover:

the equation is \[A=\frac{ a+b }{ 2 } \]

LadyRavyRay:

(11+10)(8)

ZombieLover:

you put the base 1 in (A) and put base 2 for (B)

ZombieLover:

Correct so whats the equation?

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieLover the equation is \[A=\frac{ a+b }{ 2 } \] \(\color{#0cbb34}{\text{End of Quote}}\) Forgot about the height :)

LadyRavyRay:



ZombieLover:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieLover the equation is \[A=\frac{ a+b }{ 2 } \] \(\color{#0cbb34}{\text{End of Quote}}\) Forgot about the height :) \(\color{#0cbb34}{\text{End of Quote}}\) oops well isn't it \[H \frac{ 11+10 }{ 2}\]?

LadyRavyRay:

IS the answer 84 cm2

AZ:

 Like in your screenshot \(\text{Area} = \dfrac{\text{Base}_1 + \text{Base}_2}{2} \times \text{height}\)

ZombieLover:

ohhh i was right then? the 8 is the base!

ZombieLover:

i was confused for a sec

ZombieLover:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieLover ohhh i was right then? the 8 is the base! \(\color{#0cbb34}{\text{End of Quote}}\) HIEGHT* i mean

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @LadyRavyRay IS the answer 84 cm2 \(\color{#0cbb34}{\text{End of Quote}}\) That's the area of the trapezoid Now multiply by the height of the prism which is 21

LadyRavyRay:

1,764 cm2

AZ:

Well done!

LadyRavyRay:

THANK YOU!!!

AZ:

My pleasure :D